OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = -4(x - 6)^2 + 2?

OpenStudy (zzr0ck3r):

for \(y=a(x-b)^2+c\) we have that \((b,c)\) is the vertex point.

OpenStudy (anonymous):

how do you find c

OpenStudy (anonymous):

-6,2

OpenStudy (zzr0ck3r):

\(y=a(x-b)^2+c \iff y=-4(x-6)^2+2\)

OpenStudy (zzr0ck3r):

why -6?

OpenStudy (zzr0ck3r):

notice the form has (x-b) so since you have (x-6) we have b=6

OpenStudy (anonymous):

i dont know

OpenStudy (zzr0ck3r):

if you had (x+6) then b=-6

OpenStudy (zzr0ck3r):

we want it to be \((\small x\large-\small b)\)

OpenStudy (zzr0ck3r):

so the answer is (6,2)

OpenStudy (anonymous):

ohh

OpenStudy (zzr0ck3r):

if you had \(-4(x+6)^2+2\) then the answer would be \((-6,2)\)

OpenStudy (zzr0ck3r):

because for our form to work we need it to be a minus sign, so its \(-4(x-(-6))^2+2\)

OpenStudy (zzr0ck3r):

do a few more like this and you will get what I'm saying.

OpenStudy (anonymous):

thanks